# IAFormer: Interaction-Aware Transformer network for collider data analysis

### v1

@@ 4 Analysis of hidden layer representations
We now compare the attention-based transformer models (Plain Transformer, Transformer + $\mathcal{I}_{ij}$ and IAformer) using several interpretability methods to study how hidden layers are organized for classification. Different interpretability methods can be applied to analyze attention-based Transformer models designed for particle cloud data. These techniques provide insights into how the model processes and prioritizes different particles, helping to understand learned representations and the network decision-making. By leveraging these interpretability techniques, one can gain deeper insights into the behavior of attention-based Transformers for particle cloud analysis, ensuring that their predictions align with meaningful physical patterns. Common interpretation methods are
@@
  • Saliency Maps: These maps measure the sensitivity of the model’s output to small perturbations in the input [76]. Saliency maps help assess whether the model captures physically meaningful features or relies on spurious correlations.
@@
In the following, we focus on attention maps and CKA similarity to interpret the results obtained from the Top tagging. The attention maps utilized to visualize the sparse pattern in the attention score when using IAFormer, while the CKA is used to explore the learned representation by the different attention layers.
@@ 4.1 Attention Maps
The attention map is a 2D heatmap of attention scores that highlights where the model focuses. To compute attention maps, the attention scores, which represent the strength of the relationship between different elements of the input, are computed by taking the dot product between the query vector of one element and the key vector of every other element in the input. These scores are then normalized using a softmax function, which ensures that they sum to one. The final attention map is a visual representation of these scores, showing which parts of the input the model “attends to” or focuses on most when producing its output. In IAFormer the attention map is computed as the difference between the learnable pairwise interaction matrix in equation 9.
@@
The attention maps for IAFormer show high attention scores concentrated on a few particle tokens across all attention heads, with the remaining particles significantly suppressed. In contrast, the plain Transformer distributes attention more broadly across a larger number of tokens. This shows IAFormers more efficiently accumulate relevant information from earlier tokens. Ideally, the network should focus on the cluster hadrons that form the three-prong structures of the top jet, as assigning high attention to a large number of tokens can hinder classification performance. Additionally, IAFormer attention maps exhibit a symmetric pattern, reflecting the symmetric nature of the pairwise interaction matrix. This symmetry arises even though we do not require the symmetry to weight factor $W_{1}$ and $W_{2}$.
@@ 4.2 CKA similarity
Centered Kernel Alignment (CKA) similarity, derived from kernel methods and alignment-based metrics, provides a powerful tool for assessing how closely two sets of learned representations correspond, whether across different layers within a neural network or between entirely separate models. Unlike traditional similarity measures such as Pearson correlation or Euclidean distance, which directly compare raw values, CKA evaluates how well representations align in high-dimensional feature space. This allows it to capture intricate, nonlinear relationships between features, making it particularly effective for analyzing deep learning models. One of the key challenges in studying neural network representations is that features are distributed across multiple neurons, and the number of neurons in hidden layers often exceeds the input dimension while varying across layers and models. This variability makes direct comparisons difficult. CKA addresses this issue by providing a quantitative framework for comparing learned representations, both within a single model and across different architectures.
@@
CKA values range from 00 to $1$, where higher values indicate strong similarity between learned representations. If two consecutive layers have high CKA similarity, it suggests that the second layer does not significantly enhance classification accuracy, implying that it could be removed without affecting performance. On the other hand, layers with low CKA similarity capture distinct aspects of the data, contributing to improved model performance by learning complementary features.
Figure 5: Linear CKA similarity for IAFormer (left), Transformer $+I_{i,j}$ (middle) and Plain Transformer (right) using 1000 test events from the Top jet dataset. The axes represent the attention layers in each network, while the colour bar indicates the CKA values.
We compute the output from each attention layer with dimensions $(100,32)$, where the first dimension corresponds to the number of particles and the second represents feature embeddings. To construct the Gram matrices, $M$ and $N$, we consider 1000 test events and average over the feature dimension, resulting in Gram matrices of size $(1000,1000)$. These matrices are then used to compute the CKA values for three different networks, IAFormer, Transformer with an interaction matrix, and a plain Transformer, as shown in Figure 5. For a fair comparison, all networks are trained with 12 attention layers, represented on the $X$ and $Y$ axes of each plot.
Attention layers of IAFormer exhibit lower CKA values, particularly in the first four layers, with CKA values ranging from 0.5 to 0.8. This suggests that these layers capture different patterns in the jet constituents, contributing to an overall improvement in classification performance. The later layers exhibit stronger internal similarity, with CKA values exceeding 0.8. The CKA structure indicates an efficient flow of attention from the early layers to the final ones. The middle plot presents the CKA values for the Transformer incorporating the interaction matrix. The CKA matrix exhibits a block-diagonal structure, where the first four layers demonstrate strong internal similarity but differ significantly from the remaining layers. Additionally, the last four layers encode nearly identical information, with a CKA value of 1. In contrast, the plain Transformer shows consistently high similarity across all layers, with CKA values ranging from 0.85 to 1. This uniformity likely contributes to its lower performance in the Top tagging task compared to other networks.
Overall, the Plain Transformer and Transformer+ $\mathcal{I}_{ij}$ stop improving first 6 layers, while IAFormer steadily improves for all layers, with stronger suppression of the fluctuation by larger $\beta$ layer by layer, likely building global variables along the increasing layers.
@@ 5 Conclusion
In this paper, we introduce IAFormer, an attention-based deep learning network that incorporates dynamic sparse attention to reduce network size while preserving classification performance. Furthermore, the standard attention weights are replaced with pairwise interactions between input particles. Combined with sparse attention, this modification reduces redundancy typically present in standard Transformer architectures.
To evaluate the network performance, we validate IAFormer on two benchmark tasks: top quark tagging and quark–gluon discrimination. In both cases, IAFormer achieves improved classification performance compared to the ParT network, while significantly reducing model size. Furthermore, IAFormer delivers performance competitive with state-of-the-art Lorentz-invariant networks. To further analyze the results, we employ various AI interpretability techniques to visualize the information learned by IAFormer’s hidden layers in comparison to other Transformer-based architectures. Specifically, we utilize attention maps and CKA similarity. The attention maps reveal a sparse structure in IAFormer attention weights, in contrast to the more scattered attention patterns observed in the original Transformer, where weights are more uniformly distributed across all particle tokens. This sparsity enables IAFormer to concentrate on the most relevant particles and suppress attention to less informative inputs, a behavior not exhibited by the original Transformer, which tends to assign attention broadly, even to less important tokens. Additionally, CKA analysis indicates lower similarity between IAFormer learned representations and those of other Transformer variants, highlighting its unique learning dynamics.
@@ Acknowledgments
This work is funded by grant number 22H05113, “Foundation of Machine Learning Physics”, Grant in Aid for Transformative Research Areas and 22K03626, Grant-in-Aid for Scientific Research (C). WE is funded by the ErUM-WAVE project 05D2022 “ErUM-Wave: Antizipation 3-dimensionaler Wellenfelder”, which is supported by the German Federal Ministry of Education and Research (BMBF).
Appendix A Public code
Although IAFormer is validated against the current DL networks for jet tagging problems, it is adopted for generic classification problems. Additionally, our code comprises three different networks that can be used by public users, IAFormer, Plain Transformer and Transformer with interaction matrix. In this section, we present the full details on how to use the implemented networks for different classification tasks. Implementation of the IAFormer model is publicly available at IAFormer. The repository is implemented using PyTorch Lightning [^3] https://github.com/Lightning-AI/pytorch-lightning, a high-level deep learning framework built on top of PyTorch [79]. It offers a modular, scalable, and user-friendly interface for training and evaluating deep learning models.
@@ Appendix A Public code
python main.py fit -config=configs/ia_former.yaml
The code is equipped with three different configuration files for the different networks. Hyperparamters of the best performance are saved in the checkpoints directory. To test the network
@@ A.1 Particle Graph Construction
  • feature_matrix: A 3D array of shape $[N_{\text{events}},N_{\text{particles}},F]$. Here, $F$ are the features, which incldues the particle four momenta, log-scaled transverse momentum and energy, relative kinematics, spatial distances, (Eq. 10) and particle identification variables, for the quark-gluon dataset.
  • adjacancy_matrix: A 3D array of shape $[N_{\text{events}},(N_{\text{particles}}\cdot(N_{\text{particles}}-1))/2,6]$ that contains pairwise interaction features between particles, which are listed in Eq. 11.
  • mask: A 2D boolean array of shape $[N_{\text{events}},N_{\text{particles}}]$ indicating valid (non-padded) particle entries in each event.
  • labels: A 1D array of shape $[N_{\text{events}}]$ that stores the ground-truth class labels.
@@ A.3 Networks structure
    • InteractionInputEncoding: Projects edge-level features ($I_{i,j}$) into a common embedding space using a series of Conv2D layers.
@@
  • Feature Aggregation and Output: The per-particle representations are pooled across the feature dimension, and passed to a linear classification head for binary classification.
@@
  • validation_step() and test_step(): Evaluate the model on validation and test datasets, respectively, computing accuracy, loss and the area under the ROC curve.
@@ A.4 Attention Module
The MultiHeadAttention class implements standard multi-head attention. It supports an optional interaction matrix $I_{i,j}$ that is projected using a Conv2D layer to match the shape of the attention logits. If provided, this interaction matrix is added to the dot-product attention scores before applying softmax. An optional attention mask (umask) can be used to ignore invalid or padded elements.
@@
The MultiHeadSparseAttention class implements a differential attention mechanism. The attention map is calculated entirely from the projected interaction matrix $I_{i,j}$ using a Conv2D layer. Two attention maps are extracted and combined via a learnable scalar $\beta$, parameterized through dot products between $\beta_{q}$ and $\beta_{k}$. This allows the model to softly subtract one attention map from another, enabling a sparse or selective focus on particle-particle interactions.
@@
    $\text{attn}=\text{Softmax}(I_{i,j})[:,:,0]-\beta\cdot\text{Softmax}(I_{i,j})[:,:,1]$

### v2

@@ 4 Analysis of hidden layer representations
We now compare the attention-based transformer models (Plain Transformer, Transformer + $\mathcal{I}_{ij}$(ParT), and IAFormer) using several interpretability methods to study how hidden layers are organized for classification. Different interpretability methods can be applied to analyze attention-based Transformer models designed for particle cloud data. These techniques provide insights into how the model processes and prioritizes different particles, helping to understand learned representations and the network decision-making. By leveraging these interpretability techniques, one can gain deeper insights into the behavior of attention-based Transformers, ensuring that their predictions align with meaningful physical patterns. Common interpretation methods are
@@
  • Saliency Maps: These maps measure the sensitivity of the model’s output to small perturbations in the input [34] model saliency maps help assess whether the model captures physically meaningful features or relies on spurious correlations.
@@
In the following, we focus on attention maps and CKA similarity to interpret the results obtained from the top tagging. The attention maps are utilized to visualize the sparse pattern in the attention score when using IAFormer, while the CKA is used to explore the learned representation by the different attention layers.
@@ 4.1 Attention Maps
The attention map is a 2D heatmap of attention scores that highlights the model focus. To compute attention maps, the attention scores, which represent the strength of the relationship between different elements of the input, are computed by taking the dot product between the query vector of one element and the key vector of every other element in the input. These scores are then normalized using a softmax function, which ensures that they sum to one. The final attention map is a visual representation of these scores, showing which parts of the input the model “attends to” or focuses on most when producing its output. In IAFormer, the attention map is computed as the difference between the learnable pairwise interaction matrix in equation 9.
@@
Compared with a plain Transformer model, the attention maps of IAFormer show high attention scores concentrated on a smaller number of final tokens across all attention heads, with the remaining particles significantly suppressed. This shows IAFormers more efficiently accumulate relevant information from earlier tokens. Ideally, the network should focus on the cluster of hadrons that form the three-prong structures of the top jet. If the network assigns high attention to a large number of tokens, it can hinder classification performance. Additionally, IAFormer attention maps exhibit a symmetric pattern over the diagnoal, reflecting the symmetric nature of the pairwise interaction matrix. This symmetry arises even though we do not require the symmetry to weight factors $W_{1}$ and $W_{2}$.
@@ 4.2 CKA similarity
One of the key challenges in studying neural network representations is that features are distributed across multiple neurons, and the number of neurons in hidden layers often exceeds the input dimension while varying across layers and models. This variability makes direct comparisons difficult. Centered Kernel Alignment (CKA) similarity, derived from kernel methods and alignment-based metrics, provides a powerful tool for assessing the independence of the captured information across the different layers. Traditional similarity measures, such as Pearson correlation or Euclidean distance, directly compare the raw hidden sector values. Instead of comparing the raw values, CKA evaluates how well learned “representations” in the hidden sector align in high-dimensional feature space, and therefore, it is more robust. The CKA similarity can capture intricate, nonlinear relationships between features, making it particularly effective for analyzing deep learning models.
@@
CKA values range from 0 to $1$, where higher values indicate strong similarity between learned representations. If two consecutive layers have high CKA similarity, it suggests that the second layer does not significantly enhance classification accuracy, implying that it could be removed without affecting performance. On the other hand, layers with low CKA similarity capture distinct aspects of the data, contributing to improved model performance by learning complementary features.
Figure 5: Linear CKA similarity for IAFormer (left), Transformer $\mathcal{I}_{i,j}$ (middle), and Plain Transformer (right) using 1000 test events from the top jet dataset. The axes represent the attention layers in each network, while the colour bar indicates the CKA values.
We compute the output from each attention layer with dimensions $(100,32)$, where the first dimension corresponds to the number of particles and the second represents feature embeddings. To construct the Gram matrices, $M$ and $N$, we consider 1000 test events and average over the feature dimension, resulting in Gram matrices of size $(1000,1000)$. These matrices are then used to compute the CKA values for three different networks, IAFormer, Transformer with an interaction matrix( similar to ParT), and Plain Transformer, as shown in Figure 5. For a fair comparison, all networks are trained with 12 attention layers, represented on the $X$ and $Y$ axes of each plot.
Attention layers of IAFormer exhibit lower CKA values, particularly in the first four layers, with CKA values ranging from 0.5 to 0.8. This suggests that these layers capture different patterns in the jet constituents, contributing to an overall improvement in classification performance. The later layers exhibit stronger internal similarity, with CKA values exceeding 0.8. The CKA structure indicates an efficient flow of attention from the early layers to the final ones. The middle plot presents the CKA values for the Transformer incorporating the interaction matrix. The CKA matrix exhibits a block-diagonal structure, where the first four layers demonstrate strong internal similarity but differ significantly from the remaining layers. Additionally, the last four layers encode nearly identical information, with a CKA value of 1. In contrast, Plain Transformer shows consistently high similarity across all layers, with CKA values ranging from 0.85 to 1. This uniformity likely contributes to its lower performance in the top tagging task compared to other networks.
Overall, Plain Transformer and Transformer+ $\mathcal{I}_{ij}$ stop improving the first 6 layers, while IAFormer steadily improves for all layers, with stronger suppression of the fluctuation by larger $\beta$ layer by layer, likely building global variables along the increasing layers.
@@ 5 Conclusion
In this paper, we introduce IAFormer, an attention-based deep learning network that incorporates dynamic sparse attention to reduce network size while preserving classification performance. Furthermore, the standard attention weights are replaced with pairwise interactions between input particles. This ensures the attention matrix preserves the invariance under boost along the beam direction. Combined with sparse attention, this modification reduces the redundancy typically present in standard Transformer architectures.
To evaluate the network performance, we validate IAFormer on two benchmark tasks, top quark tagging and quark–gluon discrimination. In both cases, IAFormer achieves improved classification performance compared to ParT network, while significantly reducing model size. Furthermore, IAFormer delivers performance that is competitive with state of the art Lorentz-invariant networks.
To further analyze the results, we employ various AI interpretability techniques to visualize the information learned by IAFormer hidden layers in comparison to other Transformer-based architectures. Specifically, we utilize attention maps and CKA similarity. The attention maps reveal IAFormer condenses the information efficiently in fewer tokens, in contrast to the scattered attention patterns observed in the original Transformer, where weights are more uniformly distributed across all particle tokens. This shows that the dynamic sparse attention enables IAFormer to concentrate on the most relevant tokens and suppress attention to less informative inputs. The behavior is not exhibited by the original Transformer, which tends to assign attention broadly, even to less important tokens. Additionally, CKA analysis indicates that the similarity between the learned representations among different layers of IAFormer is lower than that of other Transformer variants, highlighting its unique learning dynamics.
@@
To better understand the role of the trainable suppression parameter $\beta$ in IAFormer, we added a regularization term to the loss that discourages sharp changes in $\beta$ across layers. This helps the model learn a smoother, more structured evolution of $\beta$, guided by layer-wise scaling factors. We found that this leads to the same layer-wise $\beta$ pattern reported earlier, a rise and fall across layers, which supports its interpretation as reflecting the effective degrees of freedom needed to separate signal from background. This reinforces the idea that $\beta$ captures physically meaningful information that shapes the model’s sparse attention behavior.
@@ Acknowledgments
This work is funded by grant number 22H05113, “Foundation of Machine Learning Physics”, Grant in Aid for Transformative Research Areas and 22K03626, Grant-in-Aid for Scientific Research (C). WE is funded by the ErUM-WAVE project 05D2022, “ErUM-Wave: Antizipation 3-dimensionaler Wellenfelder”, which is supported by the German Federal Ministry of Education and Research (BMBF).
Appendix A Code Availability
Although"IAFormer is validated against the current DL networks for jet tagging problems, it is adopted for generic classification problems. Additionally, our code comprises three different networks that can be used by public users; IAFormer, Plain Transformer, and Transformer with interaction matrix(similar to ParT). In this section, we present the full details on how to use the implemented networks for different classification tasks. Implementation of the IAFormer model is publicly available at IAFormer (https://github.com/wesmail/IAFormer). The repository is implemented using PyTorch Lightning [^3] https://github.com/Lightning-AI/pytorch-lightning, a high-level deep learning framework built on top of PyTorch [67]. It offers a modular, scalable, and user-friendly interface for training and evaluating deep learning models.
@@ Appendix A Code Availability
python main.repository’snfig=configs/ia_former.yaml
The code is equipped with three different configuration files for the different networks. Hyperparameters of the best performance are saved in the checkpoints directory. To test the network
@@ A.1 Particle Graph Construction
  • feature_matrix: A 3D array of shape $N_{\text{events}},N_{\text{particles}},F]$. Here, $F$ are the features, which include the particle four momenta, log-scaled transverse momentum and energy, relative kinematics, spatial distances (Eq. 10), and particle identification variables, for the quark-gluon dataset.
  • adjacancy_matrix: A 3D array of shape $N_{\text{events}},(N_{\text{particles}}\cdot(N_{\text{particles}}-1))/2,6]$ that contains pairwise interaction features between particles, which are listed in Eq. 11.
  • mask: A 2D boolean array of shape $N_{\text{events}},N_{\text{particles}}]$ indicating valid (non-padded) particle entries in each event.
  • labels: A 1D array of shape $N_{\text{events}}]$ that stores the ground-truth class labels.
@@ A.3 Networks structure
    • InteractionInputEncoding: Projects edge-level features ($\mathcal{I}_{i,j}$) into a common embedding space using a series of Conv2D layers.
@@
  • Feature Aggregation and Output: The per-particle representations are pooled across the feature dimension and passed to a linear classification head for binary classification.
@@
  • validation_step() and test_step(): Evaluate the model on validation and test datasets, respectively, computing accuracy, loss, and the area under the ROC curve.
@@ A.4 Attention Module
The MultiHeadAttention class implements standard multi-head attention. It supports an optional interaction matrix $\mathcal{I}_{i,j}$ that is projected using a Conv2D layer to match the shape of the attention logits. If provided, this interaction matrix is added to the dot-product attention scores before applying softmax. An optional attention mask (umask) can be used to ignore invalid or padded elements. The attention mask offers the advantage of excluding padded particle tokens from the attention computation. In practice, if an event contains fewer particles than the maximum sequence length, the remaining positions are padded with zeros. Without masking, these zero-padded tokens may receive nonzero attention scores, e.g. 1, after the softmax. To address this, an attention mask is applied to assign a value of -$\infty$ to the padded tokens before the softmax operation. As a result, their attention scores become effectively zero after the softmax function, ensuring they do not influence the output.
@@
The MultiHeadSparseAttention class implements a differential attention mechanism. The attention map is calculated entirely from the projected interaction matrix $\mathcal{I}_{i,j}$ using a Conv2D layer. Two attention maps are extracted and combined via a learnable scalar $\beta$, parameterized through dot products between $\beta_{q}$ and $\beta_{k}$. This allows the model to softly subtract one attention map from another, enabling a sparse or selective focus on particle-particle interactions.
@@
    $\text{attn}=\text{Softmax}(\mathcal{I}_{i,j})[:,:,0]-\beta\cdot\text{Softmax}(\mathcal{I}_{i,j})[:,:,1]$
@@
A.5 JetClass IAFormer Implementation
This subsection describes the specific implementation of IAFormer for the JetClass benchmark, including data conversion, storage format, data loading, and model training.
JetClass Data Conversion and Storage
The JetClass dataset is provided in ROOT format, containing per-jet constituent information with variable-length particle lists. To enable efficient large-scale training, we implement a dedicated ROOT-to-HDF5 conversion pipeline. The converter reads jet constituent four-momentum components and auxiliary features, constructs particle-level node features and pairwise interaction features, and stores the result in sharded HDF5 files.
Each HDF5 file contains the following datasets:
  • node_features: A tensor of shape $N_{\text{events}},N_{\text{particles}},F]$, storing per-particle features such as kinematic variables and derived quantities.
  • adjacency_matrix: A compact upper-triangular representation of pairwise interaction features with shape $N_{\text{events}},N_{\text{pairs}},U]$, where $N_{\text{pairs}}=N_{\text{particles}}(N_{\text{particles}}-1)/2$.
  • labels: A one-dimensional array of integer class labels for each event.
  • n_particles: The true number of particles per event before padding.
Storing the interaction matrix in upper-triangular form significantly reduces disk usage and I/O overhead while preserving all pairwise information required by interaction-aware attention. Dataset-level metadata (feature definitions, maximum particle count, class mapping) are stored as HDF5 attributes for reproducibility.
Data Handling and Graph Reconstruction
The H5Dataset class in data/ provides efficient access to the JetClass HDF5 files and integrates directly with the PyTorch Lightning DataModule. During loading, the dataset:
  • infers the number of particles from n_particles,
  • reconstructs the dense or sparse particle-particle adjacency structure from the stored upper-triangular representation,
  • builds edge_index and edge_attr tensors compatible with PyTorch Geometric-style batching,
  • applies masking to padded particles to ensure they do not contribute to attention or pooling.
To support large-scale JetClass training, HDF5 file handles are cached using an LRU strategy, and static graph structures (e.g. upper-triangular indices) are cached once per process.
Model Architecture for JetClass
For JetClass classification, IAFormer operates on particle sequences with explicit interaction modeling. The model consists of:
  • Particle Embedding: An MLP that projects node-level particle features into a common embedding space.
  • Interaction Encoding: A learnable projection of pairwise interaction features, used to modulate attention scores.
  • Transformer Blocks: A stack of particle-attention blocks supporting plain, interaction-aware, or sparse differential attention.
  • Pooling and Classification: Masked pooling over particle embeddings followed by a linear classification head for multiclass jet tagging.
The choice of attention mechanism (plain, interaction, or diff) is controlled via the attn argument in the configuration files.
Training and Evaluation
Training, validation, and testing are implemented in a dedicated PyTorch Lightning module. The module computes the cross-entropy loss for multiclass classification and logs standard performance metrics, including accuracy and area under the ROC curve. During testing, attention maps and intermediate representations can optionally be stored for downstream interpretability studies.